Tuesday, November 25, 2014

Calibration Maneuvers on a Tuesday?

Today SDO did two calibration maneuvers. The EVE field of view started at 1315 UTC (8:15 a.m. ET) and the HMI/AIA flatfield started at 1630 UT (11:30 a.m. ET) and ended at 1900 UTC (2:00 p.m. ET). These motions are used by the instrument teams to track the performance of the instruments. They had been scheduled for October but were delayed to track AR 12192 across the disk and watch its re-emergence.

Here is an AIA 171 image at 1540 UTC (10:40 a.m. ET) showing a nice coronal loop in the northern hemisphere.

Friday, November 14, 2014

Another Lunar Transit, November 22, 2014

SDO will experience the third lunar transit this year on November 22, 2014. Here is the FOT movie predicting what we will see during the transit.

Wednesday, November 12, 2014

EVE Cruciform Today

The EVE cruciform is today starting at 1800 UTC. During this maneuver the spacecraft rocks back and forth, helping EVE to maintain an accurate calibration.

Tuesday, November 11, 2014

It's Coming Around Again!

Active Region 12192, which was the location of 6 X-class and several M-class flares in late October, is coming back for an encore. Here is a movie of the Sun showing the magnetic field on the side facing the Earth and the farside image from HMI data. It shows the part of the Sun containing AR 12192 rotating back into view, with a large magnetic feature still there.
These composite maps of the Sun’s farside (in yellow) and the line-of-sight magnetic field (blue) show NOAA AR12192 as a bight-dark magnetic feature rotating out of view over the west limb at the beginning, passing the far-side meridian in the middle as a dark area, and approaching the east limb at the end. The farside information is shown as the travel time change in the echo produced by a strong magnetic region.

This farside images allow us to track magnetic features where SDO cannot see. They have been compared with the full-Sun images we make by combining SDO and STEREO images. When AR 12192 returns it will have a new number but we hope it will be just as active.

Many thanks to the HMI team for producing these images and movies. They are available at JSOC.

Friday, November 7, 2014

Repairs are Complete, JSOC is Live!

The fiber optics cable is repaired and data is flowing to the JSOC. It's good to see the Sun!

Wednesday, November 5, 2014

JSOC Images Unavailable (Updated 11/6/2014)

The SDO JSOC is recovering from an internal disruption and images are unavailable. This outage should last 24 hours.

Update 11/5/2014 2142 UTC: We have learned that the fiber cables between WSMR and Stanford were accidently cut. This outage will continue until the location has been determined and the lines repaired or bypassed. The problem at the JSOC has been fixed and, until this new outage, the image cadence was being restored. 

Update 11/6/2014 1447 UTC (9:47 am ET): The fiber optics provider has found the break, dug it up, and should be splicing the fiber links soon. I do not know how long each fiber takes to splice or how many fibers are in this bundle, so there is no estimate at this time when we will return to service.